OpenStudy (anonymous):

Which is the equation to the line passing through points (2, –7) and (5, –1)? y = – 2x + 6 y = 2x – 11 y = 2x – 8 y = – 2x + 9

OpenStudy (anonymous):

to get the equation of the line passing through the points (2, –7) and (5, –1) : 1º find the slope: -1-(-7)/5-2=6/3=2 2º substitute (a,b) by the values (x,y) of any of two points into this equation: y-b=2(x-a) 3º put in the form of the equations you have in the answer set to choose the right one